# SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking

### v2

@@ 3.1 Appearance-based Retrieval
To enhance cross-modal alignment, we fine-tune the dual-encoder on these appearance-related image-text pairs from ScenePerson-13W, following the implicit relation reasoning proposed in jiang2023cross . The loss is formally expressed as:
$\mathcal{L}_{\text{app}}=\mathbb{E}_{(b_{ij},\mathcal{T}_{\text{app}},y_{ij})\sim\mathcal{D}_{\text{app}}}\left[\ell_{\text{IRR}}(f_{ij}^{v},f^{t},y_{ij};\Theta)+\ell_{\text{SDM}}(f_{ij}^{v},f^{t},y_{ij};\Theta)+\ell_{\text{ID}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\right],$ (2)
where $\mathcal{D}_{\text{app}}$ is the set of appearance-related image-text pairs with identity labels $y_{ij}$, $\ell_{\text{IRR}}$, $\ell_{\text{SDM}}$, and $\ell_{\text{ID}}$ are the respective loss components as defined in  jiang2023cross , and $\Theta$ denotes the parameters of the image and text encoders. The expectation is computed over the training set, ensuring discriminative and robust feature representations.
Retrieval. During inference, we first apply a YOLO-based person detector to a gallery of full-scene images $\mathcal{G}=\{\mathbf{I}_{j}\}_{j=1}^{M}$, extracting all detected person crops across the gallery. This yields a set of $K$ cropped person images $\mathcal{C}=\{\mathbf{C}_{k}\}_{k=1}^{K}$, where each $\mathbf{C}_{k}$ originates from a full-scene image $\mathbf{I}_{j}\in\mathcal{G}$. For each crop $\mathbf{C}_{k}$, we also retain its bounding box $\mathbf{B}_{k}$ and its source image identifier $j=s(k)$, indicating that $\mathbf{C}_{k}$ was cropped from $\mathbf{I}_{s(k)}$.
Each cropped image $C_{k}$ is encoded into $f_{k}^{v}$ using the fine-tuned image encoder and the appearance-related textual query $\mathcal{T}_{\text{app}}$ is encoded into $f^{t}$, as defined previously in Eq.(1). The similarity score between the textual feature and each person image is computed as:
@@
3.2 Scene-aware Re-ranking
While the appearance-based retrieval stage provides a strong initial filter by matching at the region-level focused on appearance, it lacks the ability to incorporate broader contextual cues, such as spatial configurations, interactions, or referential expressions. To complement this, we introduce a scene-aware ranking module based on InternVL 2.5-8B chen2024expanding , an advanced multimodal large language model (MLLM) capable of joint visual-linguistic reasoning over complex scenes and full descriptions, without any task-specific fine-tuning.
@@ 3.2 Scene-aware Re-ranking
Scene-aware Multimodal Prompt. To guide SceneRanker to perform reranking, we construct a unified prompt that integrates both the annotated visual inputs and an instructional textual query. The MLLM input is a structured multimodal prompt combining the $K$ full-scene images $\{I_{m}\}_{m=1}^{K}$, their corresponding bounding box annotations $\{\mathbf{B}_{m}\}_{m=1}^{K}$, and the complete textual query $T$. The visual prompt is formatted as a markdown sequence, where each full-scene image is paired with its bounding box annotation:
@@
[…]
Image-1: <image><box>$B_{1}$</box> Image-2: <image><box>$B_{2}$</box> Image-$K$: <image><box>$B_{K}$</box> For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7, $\cdots$, $K$].
Re-ranking. The SceneRanker processes the multimodal input ($(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P})$ to produce a ranked list of indices:
@@
4 Experiments
4.1 Experimental Setup
Dataset and Evaluation Metrics. All experiments are conducted on the proposed ScenePerson-13W dataset, which is divided into $121,438$ image-text pairs for training and $10,000$ image-text pairs for testing. Following prior retrieval works bai2023rasa ; cao2024empirical , we evaluate overall retrieval performance using Rank@k (R@k) for k=1, 5, and 10, and mean Average Precision (mAP). R@k measures the proportion of queries where the target individual is correctly identified within the top k candidates, while mAP quantifies the ranking quality across all queries.
Implementation Details. For the appearance-based retrieval stage, we adopt YOLOv5wu2021application for pedestrian detection and MediaPipe mediapipe_pose for filtering to ensure valid person instances. The dual-encoder model is fine-tuned following the original IRRA jiang2023cross implementation, initialized with CLIP-ViT-B/16 radford2021learning for images and CLIP text Transformer radford2021learning for text. The model is trained for $60$ epochs on the ScenePerson-13W dataset using the Adam optimizer with an initial learning rate of $10^{-5}$ and a cosine learning rate decay schedule. The image resolution of input crops is resized to $384\times 128$ pixels. During inference, the size of candidate set $K$ selected for further processing is set to $10$. For the SceneRanker, we use InternVL 2.5-8B chen2024expanding without task-specific fine-tuning. The generation is performed with a maximum output length of $128$ tokens and deterministic decoding. The maximum number of patches that an input image is divided into during preprocessing is set to $10$ to balance computational efficiency and detailed visual analysis. All other generation hyperparameters retain their default values as defined by the Hugging Face Transformers wolf2019huggingface library. All experiments are conducted on four RTX A40 GPUs with 48 GB of memory each.
4.2 Comparison with State-of-the-art Methods
We compare the proposed SA-Person method with a range of state-of-the-art retrieval baselines, grouped by their retrieval strategies and visual grounding scope. Region-level methods solely focus on cropped pedestrian images, emphasizing appearance cues but lacking broader scene context. Global-level methods operate on the entire scene with full textual descriptions, capturing contextual information but often sacrificing accurate grounding of pedestrian appearance. In contrast, SA-Person adopts a region-to-global retrieval paradigm, initially matching pedestrian-specific regions to ensure discriminative grounding, followed by scene-aware reranking to incorporate semantic information from the broader environment. As shown in Table 2, SA-Person consistently outperforms all baselines by a large margin. For region-level methods, it yields a significant boost of 15.41% at R@1 and 10.43% at mAP. Against the strongest region-level model, IRRA (ViT-L/14), SA-Person still improves R@1 by 14.97% and mAP by 10.15%. For global-level methods, SA-Person exceeds CLIP (ViT-L/14) by 14.83% at R@1 and 9.53 at mAP, and outperforms the best-performing LongCLIP (ViT-L/14) by 12.59% at R@1 and 7.81% at mAP. These results collectively highlight the superiority of our region-to-global cascade retrieval strategy, which integrates region-level discrimination with scene-level context and enhance both localization and semantic reasoning for more robust text-to-person retrieval.
Table 2: Comparison of the proposed SA-Person with state-of-the-art methods.
Method Focus Backbone R@1 (%) R@5 (%) R@10 (%) mAP
@@ 4.2 Comparison with State-of-the-art Methods
RaSa bai2023rasa Appearance ViT-B/16 62.51 83.83 88.09 72.05
IRRA jiang2023cross Appearance ViT-B/16 60.57 83.35 87.86 70.65
IRRA jiang2023cross Appearance ViT-L/14 63.37 84.54 88.67 72.71
TBPS-CLIP cao2024empirical Appearance ViT-B/16 50.29 75.24 81.58 61.54
RDE qin2024noisy Appearance ViTB/16 61.26 83.50 88.23 71.17
MARS ergasti2024mars Appearance ViTB/16 62.93 84.26 88.65 72.43
@@
CLIP radford2021learning Scene-aided ViT-B/16 56.37 80.73 87.16 67.22
CLIP radford2021learning Scene-aided ViT-L/14 63.51 85.53 90.28 73.33
LongCLIP zhang2024long Scene-aided ViT-B/16 58.90 82.41 88.23 69.42
LongCLIP zhang2024long Scene-aided ViT-L/14 65.75 86.22 90.95 75.05
@@
SA-Person (Ours) Scene-aided ViT-B/16 77.42 86.61 87.86 81.96
SA-Person (Ours) Scene-aided ViT-L/14 78.34 87.54 88.67 82.86
Table 3: Performance comparison of text-based person retrieval with and without the proposed SceneRanker. Red numbers indicate the absolute improvement achieved by our SceneRanker.
@@
     TBPS-CLIP cao2024empirical      50.29      75.24      81.58      61.54
@@
     IRRA(ViT-B/16) jiang2023cross      60.57      83.35      87.86      70.65
@@
     IRRA(ViT-L/14) jiang2023cross      63.37      84.54      88.67      72.71
@@
     RDE qin2024noisy      61.26      83.50      88.23      71.17
      +SceneRanker      $77.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$      $86.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.18_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
     RaSa bai2023rasa      62.51      83.83      88.09      72.05
      +SceneRanker      $77.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$      $86.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
     MARS ergasti2024mars      62.93      84.26      88.65      72.43
      +SceneRanker      $78.32_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$      $87.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
4.3 Compatibility of the Proposed SceneRanker
@@
The proposed SceneRanker is highly compatible and can be seamlessly adapted to various text-based person retrieval methods. Specifically, these methods can function as the first stage of appearance-based retrieval in the SA-Person framework. We evaluate the compatibility of SceneRanker in Table 3, where it consistently demonstrates performance improvements. Across all baselines, SceneRanker delivers over a 15% improvement in R@1 and more than a 10% increase in mAP. Notably, it achieves significant gains for TBPS-CLIP, with R@1 increasing by 21.93% (from 50.32% to 72.25%). These consistent enhancements highlight SceneRanker as a flexible and effective component that integrates seamlessly with diverse text-based person retrieval methods. By leveraging complementary contextual cues beyond standalone pedestrian appearance, SceneRanker refines candidate selection and significantly improves retrieval results.
4.4 Ablation Studies

### v3

@@ V-A Appearance-based Retrieval
To enhance cross-modal alignment, we fine-tune the dual-encoder on these appearance-related image-text pairs from ScenePerson-13W, following the implicit relation reasoning proposed in [jiang2023cross]. The loss is formally expressed as:
$\begin{split}\mathcal{L}_{\text{app}}=\mathbb{E}_{(b_{ij},\mathcal{T}_{\text{app}},y_{ij})\sim\mathcal{D}_{\text{app}}}\big[&\ell_{\text{IRR}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\\ +&\ell_{\text{SDM}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\\ +&\ell_{\text{ID}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\big],\end{split}$ (2)
where $\mathcal{D}_{\text{app}}$ is the set of appearance-related image-text pairs with identity labels $y_{ij}$, $\ell_{\text{IRR}}$, $\ell_{\text{SDM}}$, and $\ell_{\text{ID}}$ are the respective loss components as defined in [jiang2023cross], and $\Theta$ denotes the parameters of the image and text encoders. The expectation is computed over the training set, ensuring discriminative and robust feature representations.
Retrieval. During inference, we first apply the pedestrian detector to a gallery of full-scene images $\mathcal{G}=\{\mathbf{I}_{j}\}_{j=1}^{M}$, extracting all detected person crops across the gallery. This yields a set of $K$ cropped person images $\mathcal{C}=\{\mathbf{C}_{k}\}_{k=1}^{K}$, where each $\mathbf{C}_{k}$ originates from a full-scene image $\mathbf{I}_{j}\in\mathcal{G}$. For each crop $\mathbf{C}_{k}$, we also retain its bounding box $\mathbf{B}_{k}$ and its source image identifier $j=s(k)$, indicating that $\mathbf{C}_{k}$ was cropped from $\mathbf{I}_{s(k)}$.
Each cropped image $C_{k}$ is encoded into $f_{k}^{v}$ using the fine-tuned image encoder and the appearance-related text $\mathcal{T}_{\text{app}}$ is encoded into $f^{t}$, as defined previously in Eq.(1). The similarity score between the textual feature and each person image is computed as:
@@
V-B Scene-aware Re-ranking
While the appearance-based retrieval stage provides a strong initial filter by matching at the region-level focused on appearance, it can be limited in fully leveraging broader contextual cues compared to MLLMs, which offer superior holistic reasoning in full-scene settings. To leverage these complementary strengths without the overhead of gallery-wide processing, we introduce SceneRanker, a scene-aware ranking module that applies a MLLM to refine the top candidates from the appearance-based retrieval stage, without any task-specific fine-tuning. This confines expensive inference to a compact subset for scalable disambiguation.
@@ V-B Scene-aware Re-ranking
Scene-aware Multimodal Prompt. To guide SceneRanker to perform reranking, we construct a unified prompt that integrates both the annotated visual inputs and an instructional text. The MLLM input is a structured multimodal prompt combining the $K$ full-scene images $\{I_{m}\}_{m=1}^{K}$, their corresponding bounding box annotations $\{\mathbf{B}_{m}\}_{m=1}^{K}$, and the complete text $T$. The visual prompt is formatted as a markdown sequence, where each full-scene image is paired with its bounding box annotation:
@@
where $\langle\text{image}\rangle$ is a placeholder for $\mathbf{I}_{m}$, and $\langle\text{box}\rangle\mathbf{B}_{m}\langle/\text{box}\rangle$ denotes the spatial region of interest corresponding to the candidate. And an instructional prompt is appended to define the ranking task. Specifically, the final multimodal prompt $\mathbf{P}$ takes the following form:
Image-1: <image> $B_{1}</box$> Image-2: <image> $B_{2}</box$> Image-3: <image> $B_{3}</box$> Image-K: <image> $B_{K}</box$> Instruction: For the text description: “TT”, order the images based on how accurately they reflect the overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7, $\cdots$, $K$].
Re-ranking. The SceneRanker processes the multimodal input $(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P})$ to produce a ranked list of indices:
@@
VI Experiments
VI-A Experimental Setup
Dataset and Evaluation Metrics. Experiments are conducted on four benchmarks: the proposed ScenePerson-13W dataset, augmented CUHK-SYSU [xiao2017joint], Urban1K [zhang2024long], and a 5K subset of ShareGPT4V [chen2024sharegpt4v]. Unless otherwise specified, performance metrics are reported on ScenePerson-13W. Following prior retrieval works [bai2023rasa, cao2024empirical], we evaluate overall retrieval performance using Rank@k (R@k) for k=1, 5, and 10, respectively, and mean Average Precision (mAP). R@k measures the proportion of queries where the target individual is correctly identified within the top k candidates, while mAP quantifies the ranking quality across all queries.
TABLE III: Comparison of the proposed SA-Person with state-of-the-art methods.
Method Venue Focus Backbone R@1 (%) R@5 (%) R@10 (%) mAP
@@ VI-A Experimental Setup
RaSa [bai2023rasa] IJCAI’23 Appearance ViT-B/16 62.51 83.83 88.09 72.05
IRRA [jiang2023cross] CVPR’23 Appearance ViT-B/16 60.57 83.35 87.86 70.65
IRRA [jiang2023cross] CVPR’23 Appearance ViT-L/14 63.37 84.54 88.67 72.71
TBPS-CLIP [cao2024empirical] AAAI’24 Appearance ViT-B/16 50.29 75.24 81.58 61.54
RDE [qin2024noisy] CVPR’24 Appearance ViTB/16 61.26 83.50 88.23 71.17
MARS [ergasti2024mars] TOMM’24 Appearance ViTB/16 62.93 84.26 88.65 72.43
FMFA [yin2025cross] arXiv’25 Appearance ViTB/16 59.77 83.22 88.05 70.09
@@
CLIP [radford2021learning] ICML’21 Scene-aided ViT-B/16 56.37 80.73 87.16 67.22
CLIP [radford2021learning] ICML’21 Scene-aided ViT-L/14 63.51 85.53 90.28 73.33
LongCLIP [zhang2024long] ECCV’24 Scene-aided ViT-B/16 58.90 82.41 88.23 69.42
LongCLIP [zhang2024long] ECCV’24 Scene-aided ViT-L/14 65.75 86.22 90.95 75.05
@@
SA-Person Ours Scene-aided ViT-B/16 77.42 86.61 87.86 81.96
SA-Person Ours Scene-aided ViT-L/14 78.34 87.54 88.67 82.86
Implementation Details.
For the appearance-based retrieval stage, the pedestrian detector is composed of YOLOv5 [wu2021application] for initial detection and MediaPipe [mediapipe_pose] for keypoint filtering, ensuring valid and reliable person instances. The dual-encoder model is fine-tuned following the original IRRA [jiang2023cross] implementation, initialized with CLIP-ViT-B/16 [radford2021learning] for images and CLIP text Transformer [radford2021learning] for text. The model is trained for $60$ epochs using the Adam optimizer with an initial learning rate of $10^{-5}$ and a cosine learning rate decay schedule. The image resolution of input crops is resized to $384\times 128$ pixels. During inference, the size of candidate set $K$ selected for further processing is set to $10$. For the scene-aware re-ranking stage, we employ InternVL 2.5-8B [chen2024expanding] as SceneRanker, without task-specific fine-tuning. Inference is performed with a maximum of $128$ output tokens and deterministic decoding. All other settings follow the default Hugging Face Transformers [wolf2019huggingface] configuration. All experiments are conducted on four RTX A40 GPUs with 48 GB of memory each.
VI-B Comparison with State-of-the-art Methods
We compare the proposed SA-Person with a range of state-of-the-art retrieval baselines, grouped by their retrieval strategies and visual grounding scope. Region-level methods solely focus on cropped pedestrian images, emphasizing appearance cues but lacking broader scene context. Global-level methods operate on the entire scene with full textual descriptions, capturing contextual information but often sacrificing accurate grounding of pedestrian appearance. In contrast, SA-Person adopts a region-to-global retrieval paradigm, initially matching pedestrian-specific regions to ensure discriminative grounding, followed by scene-aware reranking to incorporate semantic information from the broader environment. As shown in Table III, SA-Person consistently outperforms all baselines by a large margin. For region-level methods, SA-Person yields a significant boost of 14.97% at R@1 and 10.15% at mAP over the strongest method IRRA (ViT-L/14). For global-level methods, SA-Person exceeds CLIP (ViT-L/14) by 14.83% at R@1 and 9.53 at mAP, and outperforms the best-performing LongCLIP (ViT-L/14) by 12.59% at R@1 and 7.81% at mAP. These results collectively highlight the superiority of our region-to-global cascade retrieval strategy, which integrates region-level discrimination with scene-level context and enhance both localization and semantic reasoning for more robust text-to-person retrieval.
TABLE IV: Performance comparison of text-based person retrieval methods with and without the proposed SceneRanker. Red numbers indicate the absolute improvement achieved by our SceneRanker.
@@ VI-B Comparison with State-of-the-art Methods
TBPS-CLIP [cao2024empirical] 50.29 75.24 81.58 61.54
@@
IRRA(ViT-B/16) [jiang2023cross] 60.57 83.35 87.86 70.65
@@
IRRA(ViT-L/14) [jiang2023cross] 63.37 84.54 88.67 72.71
@@
RDE [qin2024noisy] 61.26 83.50 88.23 71.17
+SceneRanker $77.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$ $6.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}} 88$$8.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.18_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
RaSa [bai2023rasa] 62.51 83.83 88.09 72.05
+SceneRanker $77.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$ $6.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}} 88$$8.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
MARS [ergasti2024mars] 62.93 84.26 88.65 72.43
+SceneRanker $78.32_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$ $7.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}} 88$$8.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
FMFA [yin2025cross] 59.77 83.22 88.05 70.09
+SceneRanker $77.35_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+17.58}}$ $6.59_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.37}} 88$$8.05_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.95_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.86}}$
VI-C Compatibility of the Proposed SceneRanker
@@
The proposed SceneRanker, exploiting global scene context for precise disambiguation in the scene-aware reranking stage, is highly compatible and can be seamlessly adapted to various text-based person retrieval methods. Specifically, these methods can function as the first stage of appearance-based retrieval in the SA-Person framework. We evaluate the compatibility of SceneRanker in Table IV, where it consistently demonstrates performance improvements. Across all baselines, SceneRanker delivers over a 14% improvement in R@1 and more than a 10% increase in mAP. Notably, it achieves significant gains for TBPS-CLIP, with R@1 increasing by 21.96% (from 50.29% to 72.25%). These consistent enhancements highlight SceneRanker as a flexible and effective component that integrates seamlessly with diverse text-based person retrieval methods. By leveraging complementary contextual cues beyond standalone pedestrian appearance, SceneRanker refines candidate selection and significantly improves retrieval results.
VI-D Ablation Studies